# When Are Combinations of Humans and AI Useful?

### v1

@@ 4.2.1 Calculation of Effect Size
We computed Hedges’ $g$ to measure the effect of combining human and artificial intelligence on task performance [hedges1981distribution]. For strong synergy, Hedges’ $g$ represents the standardized mean difference between the performance of the human-AI system performance and the baseline, which can be the human alone or AI alone, whichever one performs better, on average. For weak synergy (better than humans alone), Hedges’ $g$ represents the standardized mean difference between the performance of the human-AI system and the baseline of the human alone.
@@ 4.3 Bias Tests
Figure S7 displays the funnel plot of the included effect sizes, and we do not observe significant asymmetry or regions of missing data in the plot for our primary outcome, strong synergy. Egger’s regression did not indicate evidence of publication bias in the sample ($\beta=-0.67,p=0.4378$), nor did the rank correlation test ($\tau=0.05,p=0.1211$). Taken as a whole, these tests suggest that our results for strong synergy are robust to potential publication bias.
Importantly, however, we do find potential evidence of publication bias in favor of studies that report results in which the human-AI system outperforms the human alone (weak synergy). In this case, Egger’s regression does point to publication bias in the sample ($\beta=1.95,p=0.0016$), as does the rank correlation test ($\tau=0.19,p<0.0001$).
The discrepancy between the symmetry in the funnel plot for strong synergy versus asymmetry in the funnel plot for weak synergy may reflect how many researchers and journals implicitly assume an interest in weak synergy, comparing the human-AI system to the human alone.
@@ 4.4 Sensitivity Analysis
As a robustness check, though, we performed a sensitivity re-analysis in which we clustered at the paper level instead of the experiment level. This multilevel model accounts for variance in the observed effect sizes (first level), variance between effect sizes from the same paper (second level), and variance between papers (third level). We still calculated cluster-robust standard errors, confidence intervals, and statistical tests for our effect size estimates in which we defined clusters at the level of the experiment because the participant samples only overlapped on the level of the experiment. Using this approach, we found a comparable overall effect size for strong synergy of $g=-0.25$ ($p=0.0008$) and for weak synergy of $g=0.60$ ($p<0.0001$).
We also evaluated the robustness of our results to outlying and influential data points. To detect such data, we computed the residuals and Cook’s distance for each effect size. We considered residual values greater or less than three standard deviations from the mean as outliers, and following [altman2016analyzing], we considered values greater than $4/N$ as high influence, where $N$ is the number of data points in our analysis. Using this approach, we identified 11 outliers. We performed a sensitivity re-analysis on a dataset excluding these effect sizes, which resulted in similar effect size for strong synergy ($g=-0.24$, $p=0.0010$, 95% CI $[-0.38,-0.10]$).
Additionally, we conducted leave-one-out analyses, in which we performed a series of sensitivity re-analysis on the different subsets of effect sizes obtained by leaving out one effect size in our original dataset. We also conducted leave-one-out analyses at the experiment and publication levels. These tests show how each effect size, experiment, and publication affect our overall estimate of the effect of human-AI collaboration on task performance. Summary effect sizes for strong synergy ranged from $-0.28$ to $-0.19$ with $p<0.05$ in all cases, indicating the robustness of our results to any single effect size, experiment, or paper; likewise, summary effect sizes for weak synergy (better than humans alone) ranged from $0.61$ to $0.66$ with $p<0.05$ in all cases, indicating the robustness of our results to any single effect size, experiment, or paper.
Lastly, we conducted a sensitivity re-analysis on a dataset that omits the effect sizes we estimated, either using WebPlotDigitizer or the information provided by the authors in their paper, and again we found almost identical results ($g=-0.18$, $p=0.0254$, 95% CI $[-0.34,-0.02]$).
@@
We compiled the data used in this analysis based on the studies identified in our systematic literature review. We make the data we collected available via the Open Science Framework repository (https://osf.io/wrq7c/?view_only=b9e1e86079c048b4bfb03bee6966e560).
@@ 5 Acknowledgements
MV was funded by the Accenture Technology Convergence Fellowship. Additional funding was provided by Toyota Research Institute, by the MIT Quest for Intelligence, and by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) program. The funding organizations had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript. We also thank Jessika Kim and Garrett Campagna for their help collecting the data for this project.
6 Author Contributions
@@ 7 Competing Interests
S1 Types of Human-AI Synergy
@@
Definition 1 (Strong Synergy).
@@
Definition 2 (Weak Synergy: Better than Human Alone).
@@
Definition 3 (Weak Synergy: Better than AI Alone).
@@
S2 Details on Methods
S2.1 Systematic Literature Review
Table S1: Syntax of the search strings for the literature review.
@@
S2.2 Data Collection and Coding
@@
S2.3 Calculation of Effect Size
@@
S3 Additional Results
S3.1 Descriptive Statistics
@@
Table S3: Results from the three-level meta-regression models for the moderator variables. Here, $K$ is the number of included effect sizes for the moderator subgroup level, $g$ is the estimated effect size with the corresponding 95% confidence interval for the subgroup level, and $p$ is the $p$-value of the estimated effect size for the subgroup level.
@@
Figure S3: Forest plots of all effect sizes ($k=367$) included in the meta-analysis for weak (AI) synergy and negative synergy. The positions of the points on the $x$-axes represent the values of the effect sizes, and the bars indicate the 95% confidence interval for the effect sizes. The colors of the points and lines correspond to the values of the effect sizes, with negative effect sizes (no human-AI synergy) colored red and positive effect sizes colored green (human-AI synergy). The black dotted line corresponds to an effect size of Hedges’ $g=0$, which means that the human-AI group performed the same as the baseline. The circle at the bottom of the graph represents the meta-analytic average effect size and confidence interval.
@@
Figure S4: Scatterplot of the effect sizes that correspond to decision tasks where the researchers measured performance according to accuracy. The position of the points on the $x$-axis represents the accuracy of the human alone and the position of the points on the $y$-axis represents the accuracy of the AI alone. The colors of the points correspond to the value of the effect sizes, with negative effect sizes red and positive effect sizes green. The gray dotted line indicates that the accuracy of the humans alone and AI alone are equivalent. In the region above this line, the accuracy of the AI alone exceeds that of the human alone, and we observe a large proportion of red points – no evidence of strong human-AI synergy. In the region below this line, the accuracy of the humans alone exceeds that of the AI alone, and we observe a large proportion of green points – evidence strong human-AI synergy.
@@
S3.2 Heterogeneity Analysis
Measure Strong Synergy Weak Synergy*
$\tau^{2}$ 1.19 0.40
$I^{2}$ 97.7% 93.8%
$I^{2}_{(2)}$ 23.9% 43.6%
$I^{2}_{(3)}$ 73.8% 50.2%
$Q$-Test $p<0.0001$ $p<0.0001$
Table S4: Heterogeneity Analysis. *Better than human alone.
S3.3 Division of Labor
Only 3 of the 100+ experiments in our analysis explore processes with a pre-determined delegation of separate sub-tasks to humans and AI, so the difference between effect sizes with and without division of labor was not statistically significant. However, with the 4 effect sizes from these 3 experiments, we did find that positive human-AI synergy ($g=0.22,N.S.$) occurred among experiments with a predetermined division of labor between the human and AI, while in the experiments without this feature, the effect size was significantly negative ($g=-0.21,p=0.0102$). For example, in Lee et al. (2021), physical therapists review patient-specific analysis and then provide feature-based feedback to an AI-system, which then generates an assessment for that patient based on its training data as well as this new information [lee2021a]. Here, the authors find that their human-AI system achieves greater accuracy (93%) than either the therapist alone (75%) or AI alone without the therapist input (87%).
@@
Lastly, Lai et al. (2022) also create a pre-determined division of labor where an AI system generates a summary of a post, which a human can edit and refine [lai2022exploration]. They find, however, that the human alone produces higher quality summaries (5.8/7) than both the AI alone (4.7/7) and the human-AI combination (5.5/7). So this particular task allocation does not lead to human-computer synergy.
Figure S7: Funnel plots of the observed effect sizes and corresponding standard errors for strong and weak human-AI synergy. The blue dots represent effect sizes and the vertical dotted black line represents the estimated pooled effect sizes from the meta-analytic models. The area shaded in white represents regions of statistical non-significance ($p>0.05$), and the area shaded in gray represents regions of statistical significance ($p\leq 0.05$). In the absence of publication bias, we expect the points to fall roughly symmetrically around the estimated pooled effect size with a large proportion within the white triangle. We observe this pattern in the plot for strong human-AI synergy, but less so in the plots for weak human-AI synergy.

### v2

@@ 4.2.1 Calculation of Effect Size
We computed Hedges’ $g$ to measure the effect of combining human and artificial intelligence on task performance [hedges1981distribution]. For strong synergy, Hedges’ $g$ represents the standardized mean difference between the performance of the human-AI system performance and the baseline, which can be the human alone or AI alone, whichever one performs better, on average. For human augmentation, Hedges’ $g$ represents the standardized mean difference between the performance of the human-AI system and the baseline of the human alone.
@@ 4.3 Bias Tests
Figure S9 displays the funnel plot of the included effect sizes, and we do not observe significant asymmetry or regions of missing data in the plot for our primary outcome, human-AI synergy. Egger’s regression did not indicate evidence of publication bias in the sample ($\beta=-0.67$, $t(104)=-0.78$, two-tailed $p=0.438$, 95% CI $-2.39$ to $1.04$), nor did the rank correlation test ($\tau=0.05$,two-tailed $p=0.121$). Taken as a whole, these tests suggest that our results for human-AI synergy are robust to potential publication bias.
Importantly, however, we do find potential evidence of publication bias in favor of studies that report results in which the human-AI system outperforms the human alone (human augmentation). In this case, Egger’s regression does point to publication bias in the sample ($\beta=1.96$, $t(104)=3.24$, two-tailed $p=0.002$, 95% CI $0.76$ to $3.16$), as does the rank correlation test ($\tau=0.19$,two-tailed $p=0.000$). Note that we did not try to correct for potential publication bias to preserve the integrity of the original data and maintain transparency in our reporting. Many proposed adjustment methods can also lead to overcorrection and distort results [rothstein2005publication].
The discrepancy between the symmetry in the funnel plot for human-AI synergy versus asymmetry in the funnel plot for human augmentation may reflect how many researchers and journals implicitly assume an interest in human augmentation, comparing the human-AI system to the human alone.
@@ 4.4 Sensitivity Analysis
As a robustness check, though, we performed a sensitivity re-analysis in which we clustered at the paper level instead of the experiment level. This multilevel model accounts for variance in the observed effect sizes (first level), variance between effect sizes from the same paper (second level), and variance between papers (third level). We still calculated cluster-robust standard errors, confidence intervals, and statistical tests for our effect size estimates in which we defined clusters at the level of the experiment because the participant samples only overlapped on the level of the experiment. Using this approach, we found a comparable overall effect size for human-AI synergy ($g=-0.22$, $t(67)=-2.46$, two-tailed $p=0.017$, 95% CI $-0.41$ to $-0.04$) and for human augmentation ($g=0.65$, $t(69)=9.96$, two-tailed $p=0.000$, 95% CI $0.52$ to $0.78$).
We also evaluated the robustness of our results to outlying and influential data points. To detect such data, we computed the residuals and Cook’s distance for each effect size. We considered residual values greater or less than three standard deviations from the mean as outliers, and following [altman2016analyzing], we considered values greater than $4/N$ as high influence, where $N$ is the number of data points in our analysis. Using this approach, we identified 11 outliers for human-AI synergy and 9 outliers for human augmentation. We performed a sensitivity re-analysis on a dataset excluding these effect sizes, which resulted in similar effect size for human-AI synergy ($g=-0.25$, $t(104)=-3.45$, two-tailed $p=0.001$, 95% CI $[-0.39,-0.11]$) and human augmentation ($g=0.60$, $t(104)=12.60$, two-tailed $p=0.000$, 95% CI $[0.50,0.69]$).
Additionally, we conducted leave-one-out analyses, in which we performed a series of sensitivity re-analysis on the different subsets of effect sizes obtained by leaving out one effect size in our original dataset. We also conducted leave-one-out analyses at the experiment and publication levels. These tests show how each effect size, experiment, and publication affect our overall estimate of the effect of human-AI collaboration on task performance. Summary effect sizes for human-AI synergy ranged from $-0.28$ to $-0.19$ with two-tailed $p<0.05$ ($0.000$ to $0.019$) in all cases, indicating the robustness of our results to any single effect size, experiment, or paper; likewise, summary effect sizes for human augmentation ranged from $0.61$ to $0.66$ with two-tailed $p<0.05$ ($0.000$ to $0.000$) in all cases, indicating the robustness of our results to any single effect size, experiment, or paper.
Lastly, we conducted a sensitivity re-analysis on a dataset that omits the effect sizes we estimated, either using WebPlotDigitizer or the information provided by the authors in their paper, and again we found almost identical results for human-AI synergy ($g=-0.21$, $t(98)=-2.59$, two-tailed $p=0.011$, 95% CI $[-0.36,-0.05]$) and human augmentation ($g=0.64$, $t(98)=11.73$, two-tailed $p=0.000$, 95% CI $[0.53,0.75]$).
@@
5 Data Availability
We compiled the data used in this analysis based on the studies identified in our systematic literature review. We make the data we collected available via the project’s Open Science Framework repository (https://osf.io/wrq7c/?view_only=b9e1e86079c048b4bfb03bee6966e560). In our systematic review, we searched the following databases: the ACM Digital Library (ACM DL) (https://dl.acm.org/), Web of Science (https://clarivate.com/webofsciencegroup/solutions/web-of-science/), and Association for Information Systems eLibrary (AISeL)(https://aisnet.org/page/AISeLibrary).
6 Code Availability
@@ 7 Acknowledgements
MV was funded by the Accenture Technology Convergence Fellowship. Additional funding to TM was provided by the Toyota Research Institute, by the MIT Quest for Intelligence, and by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) program. The funding organizations had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript. We thank Jessika Kim and Garrett Campagna for their help collecting the data for this project, and we thank Dean Eckles, Mohammed Alsobay, Robin Na, and Emily Hu for their helpful feedback.
8 Author Contribution Statement
@@ 9 Competing Interest Statement
10 Figures
Figure 1: Forest plots of all effect sizes ($k=370$) included in the meta-analysis. The positions of the points on the $x$-axes represent the values of the effect sizes, and the bars indicate the 95% confidence interval for the effect sizes. The colors of the points and lines correspond to the values of the effect sizes, with negative effect sizes colored red and positive effect sizes colored green. The black dotted line corresponds to an effect size of Hedges’ $g=0$, which means that the human-AI system performed the same as the baseline. The circle at the bottom of the graph represents the meta-analytic average effect size and confidence interval.
Figure 2: Results from the three-level meta-regression models for the moderator variables. Here, $N$ is the number of included effect sizes for the moderator subgroup level, and the estimated effect size with the corresponding 95% confidence interval. The symbols in front of the moderator indicate if there is a statistically significant difference between the subgroups for human-AI synergy (*) and human augmentation ().
@@ 10 Figures
S1 Supplementary Methods
S1.1 Types of Outcomes in Human-AI Systems
@@
Definition 1 (Human-AI Synergy).
@@
Definition 2 (Human Augmentation).
@@
Definition 3 (AI Augmentation).
@@
S1.2 Systematic Literature Review
@@
Table S1: Syntax of the search strings for the literature review.
@@
S1.3 Data Collection and Coding
@@
S1.4 Calculation of Effect Size
@@
S2 Supplementary Results
S2.1 Descriptive Statistics
@@
S2.2 Effect Sizes for Additional Outcomes
Measure Human-AI Synergy Human Augmentation AI Augmentation Negative Synergy
$n$ 370 370 370 370
Hedges’ $g$0.23 0.64 0.30 1.14
95% CI [-0.39, -0.07] [0.53, 0.74] [-0.03, 0.62] [0.90, 1.38]
$t$-statistic $t(92.39)=-2.89$ $t(98.31)=11.88$ $t(99.19)=1.82$ $t(101.12)=9.58$
$p$ 0.005 0.000 0.072 0.000
Table S3: We fit separate meta-analytic models on the full set of results, and we report the results for human-AI synergy, human augmentation, AI augmentation, and negative synergy in this table. Here, $n$ corresponds to the number of effect sizes in the model, Hedges’ $g$ is the meta-analytic average effect size, 95% CI refers to the 95% confidence interval for the effect size, $t$-statistic is the test statistic with the degrees of freedom estimated according to the Satterthwaite approximation, and $p$ is the two-sided $p$-value that indicates if the effect size is statistically significant from zero.
Human-AI Synergy Human Augmentation
Better Alone AI Human AI Human
$n$ 249 121 249 121
Hedges’ $g$0.53 0.39 0.73 0.39
95% CI [-0.71, -0.35] [0.21, 0.58] [0.62, 0.84] [0.20, 0.58]
$t$-statistic $t(69.73)=-5.79$ $t(36.33)=4.27$ $t(79.38)=12.81$ $t(36.39)=4.23$
$p$ $0.000$ $0.000$ $0.000$ $0.000$
Table S4: We fit separate meta-analytic models on the subset of results where (1) the AI performs better alone and (2) the human performs better alone, and we report the results for human-AI synergy and human augmentation in this table. Here, $n$ corresponds to the number of effect sizes in the model, Hedges’ $g$ is the meta-analytic average effect size, 95% CI refers to the 95% confidence interval for the effect size, $t$-statistic is the test statistic with the degrees of freedom estimated according to the Satterthwaite approximation, and $p$ is the two-sided $p$-value that indicates if the effect size is statistically significant from zero.
@@
Figure S3: Forest plots of all effect sizes ($n=370$) included in the meta-analysis for AI augmentation and negative synergy. The positions of the points on the $x$-axes represent the values of the effect sizes, and the bars indicate the 95% confidence interval for the effect sizes. The colors of the points and lines correspond to the values of the effect sizes, with negative effect sizes (no human-AI synergy) colored red and positive effect sizes colored green (human-AI synergy). The black dotted line corresponds to an effect size of Hedges’ $g=0$, which means that the human-AI group performed the same as the baseline. The circle at the bottom of the graph represents the meta-analytic average effect size and confidence interval.
S2.3 Heterogeneity Analyses
Measure Human-AI Synergy Human Augmentation AI Augmentation Negative Synergy
$n$ 370 370 370 370
$\tau^{2}$ 1.19 0.40 4.25 2.02
$I^{2}$ 97.7% 93.8% 99.3% 98.6%
$I^{2}_{(2)}$ 23.9% 43.6% 40.7% 50.1%
$I^{2}_{(3)}$ 73.8% 50.2% 58.6% 48.5%
$Q$-Statistic $Q(369)=8532$ $Q(369)=4226$ $Q(369)=16017$ $Q(369)=7209$
$p$ $0.000$ $0.000$ $0.000$ $0.000$
Table S5: We fit separate meta-analytic models on the full set of results, and we report the heterogeneity metrics for the models of human-AI synergy, human augmentation, AI augmentation, and negative synergy. In particular, we provide $n$ (number of effect sizes), $\tau^{2}$, $I^{2}$, $I^{2}_{(2)}$, $I^{2}_{(3)}$, $Q$-statistic (with the corresponding degrees of freedom), and $p$ is the two-sided $p$-value from the $Q$-Test.
Human-AI Synergy Human Augmentation
Better Alone AI Human AI Human
$n$ 249 121 249 121
$\tau^{2}$ 1.25 0.49 0.31 0.49
$I^{2}$ 98.1% 92.7% 93.0% 93.2%
$I^{2}_{(2)}$ 17.9% 30.2% 52.5% 30.7%
$I^{2}_{(3)}$ 80.2% 62.5% 40.5% 62.6%
$Q$-statistic $Q(248)=6111$ $Q(120)=1071$ $Q(248)=2741$ $Q(120)=1136$
$p$ $0.000$ $0.000$ $0.000$ $0.000$
Table S6: We fit separate meta-analytic models on the subset of results where (1) the AI performs better alone and (2) the human performs better alone, and we report the heterogeneity metrics for the models of human-AI synergy and human augmentation. In particular, we provide $n$ (number of effect sizes), $\tau^{2}$, $I^{2}$, $I^{2}_{(2)}$, $I^{2}_{(3)}$, $Q$-statistic (with the corresponding degrees of freedom), and $p$ is the two-sided $p$-value from the $Q$-Test.
S2.4 Moderators for Additional Outcomes
Table S7: Results from the three-level meta-regression models for the moderator variables. Here, $n$ is the number of included effect sizes for the moderator subgroup level, $g$ is the estimated effect size with the corresponding 95% confidence interval for the subgroup level, and $p$ is the two-sided $p$-value of the estimated effect size for the subgroup level, which tests if the effect size is significantly different from zero.
S2.5 Scatterplots of Effect Sizes for Accuracy-Based Tasks
@@
Figure S4: Scatterplot of the effect sizes that correspond to decision tasks where the researchers measured performance according to accuracy. The position of the points on the $x$-axis represents the accuracy of the human alone and the position of the points on the $y$-axis represents the accuracy of the AI alone. The colors of the points correspond to the value of the effect sizes, with negative effect sizes red and positive effect sizes green. The gray dotted line indicates that the accuracy of the humans alone and AI alone are equivalent. In the region above this line, the accuracy of the AI alone exceeds that of the human alone, and we observe a large proportion of red points – no evidence of human-AI synergy. In the region below this line, the accuracy of the humans alone exceeds that of the AI alone, and we observe a large proportion of green points – evidence human-AI synergy.
@@
S2.6 Division of Labor
Only 3 of the 100+ experiments in our analysis explore processes with a pre-determined delegation of separate sub-tasks to humans and AI, so the difference between effect sizes with and without division of labor was not statistically significant. However, with the 4 effect sizes from these 3 experiments, we did find that positive human-AI synergy ($g=0.22$, $t(104)=0.69$, two-tailed $p=0.494$, 95% CI $-0.42$ to $0.87$) occurred among experiments with a predetermined division of labor between the human and AI, while in the experiments without this feature, the effect size was significantly negative ($g=-0.24$, $t(104)=-2.93$, two-tailed $p=0.004$, 95% CI $-0.40$ to $-0.08$). For example, in Lee et al. (2021), physical therapists review patient-specific analysis and then provide feature-based feedback to an AI-system, which then generates an assessment for that patient based on its training data as well as this new information [lee2021a]. Here, the authors find that their human-AI system achieves greater accuracy (93%) than either the therapist alone (75%) or AI alone without the therapist input (87%).
@@
Lastly, Lai et al. (2022) also create a pre-determined division of labor where an AI system generates a summary of a post, which a human can edit and refine [lai2022exploration]. They find, however, that the human alone produces higher quality summaries (5.8/7) than both the AI alone (4.7/7) and the human-AI combination (5.5/7). So this particular task allocation does not lead to human-AI synergy.
S2.7 Effect Sizes over Time
When we look at the evolution of human-AI synergy over the past years (see Figure S8), we observe potential signs of progress. This suggestive trend may reflect increased attention to designing studies that elicit human-AI synergy. Notably, however, the progress—if it exists—is not by any means rapid, and stands in stark contrast to the quick pace of development of AI systems themselves, particularly large language models (LLMs). We hope that our work can spur further progress in this domain.
Figure S8: Plots of the meta-analytic average effect sizes for human-AI synergy and human augmentation by year of publication. The error bars correspond to 95% confidence intervals, and the dotted black line corresponds to an effect size of zero (no effect).
S2.8 Bias Tests
Figure S9: Funnel plots of the observed effect sizes and corresponding standard errors for human-AI synergy, human augmentation, AI augmentation, and negative synergy. The blue dots represent effect sizes and the vertical dotted black line represents the estimated pooled effect sizes from the meta-analytic models. The area shaded in white represents regions of statistical non-significance ($p>0.05$), and the area shaded in gray represents regions of statistical significance ($p\leq 0.05$). In the absence of publication bias, we expect the points to fall roughly symmetrically around the estimated pooled effect size with a large proportion within the white triangle. We observe this pattern in the plot for human-AI synergy, but less so in the plots for human augmentation, AI augmentation, and negative synergy.